Questions Assignment Score: 48.3% Collapse Resources Hint Check Answer > ? Assignment lnfo Question 8 of 12 > Tools Suppose an oil company is thinking of buying some land for $950,000. There is a 60% chance of economic growth and a 40% chance of recession. The probability of discovering oil is 48% when there is economic growth and 34% when there is a recession. If there is cconomic growth and the oil company discovers oil, the value of the land will triplc. If they do not discover oil, the value of the land will decrease by 12%. If there is a recession and the company discovers oil, the value of the land will increase by 50%. If they do not discover oil, the land will decrease in value by 70%. What is the expected valuc of the investment? Give your answer to the ncarest dollar. Avoid rounding within calculations. Select the correct interpretation of the expected value O O O O O The expected value provides the total cost of investing in the land. The oil company should not invest in the land bccausc thc investment costs will cxcccd $950,000. The expected value is the mean investment value for the land purchase of $950,000. The oil company should not invest in thc land because the expected valuc indicates an average loss in investment. The expected value represents what the actual investment value will be for this land purchase of $950,000. The company should make the investment because the expected value predicts a profitable investment value. The expected value represents the average profit expected for a land purchase of $950,000. Thus, the company should invest in the land because the expected value is greater than zero. The expected value represents the mean investment valuc for a $950,000 land purchase. The oil company should invest in the land because, on average, the investment is profitable.
Answer #1

P( economic growth and oil discovery ) = 0.6*0.48 = 0.288, Value of investment in this case is given as: 3*950,000 = 2850000

P( economic growth and no oil ) = 0.6*0.52 = 0.312, Value of investment in this case would be (1-0.12)*950,000 = 836000

P( recession and oil discovery ) = 0.4*0.34 = 0.136, Value of investment here: 950,000*1.5 = 1425000

P( recession and no oil ) = 0.4*0.66 = 0.264, Value of investment here: 0.3*950,000 = 285000

The expected value of the investment here is computed as the sum product of the value of investment in various cases with their respective probabilities. This is computed as:

=  2850000*0.288 + 836000*0.312 + 1425000*0.136 + 285000*0.264

= 1350672

Therefore the expected value of the investment here is 1,350,672 - 950,000 = 400,672 > 0

The correct interpretation of this result is that it represents a average profit value for a 950,000 land purchase. As the expected value is more than 0, therefore the investment should be made. Therefore D is the correct answer here.